# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

### v2

@@ Abstract
Algorithmic scatter dimension is a notion of metric spaces, introduced recently by Abbasi et al. (FOCS 2023), which unifies various well-known metric spaces, including continuous Euclidean space, bounded doubling space, planar and bounded treewidth metrics. Recently, Bourneuf and Pilipczuk (SODA 2025) showed that metrics induced by graphs from any fixed proper minor closed graph class have bounded scatter dimension. Abbasi et al. presented a unified approach to obtain EPASes (i.e., $(1+\varepsilon)$-approximations running in time FPT in $k$ and $\varepsilon$) for $k$-Clustering in metrics of bounded scatter dimension. However, a seemingly inherent limitation of their approach was that it could only handle clustering objectives where each point was assigned to the closest chosen center, i.e., Voronoi clustering. They explicitly asked, if there exist EPASes for constrained $k$-Clustering for metrics of bounded scatter dimension.
We present a unified framework which yields EPASes for metrics of bounded algorithmic scatter dimension, for capacitated and fair $k$-Median/Means. Our framework exploits coresets for such constrained clustering problems in a novel manner, and notably, requires only coresets of size $\left(k\log n\varepsilon^{-1}\right)^{O(1)}$, which are usually constuctible even in general metrics. Furthermore, we note that due to existing lower bounds it is impossible to obtain such an EPAS for Capacitated $k$-Center, thereby, essentially answering the complete spectrum of the question.
Our results on capacitated and fair $k$-Median/Means make significant progress, and provide the first EPASes for these problems in broad families of metric spaces. Earlier such results were only known in continuous Euclidean spaces due to the work of Cohen-Addad & Li, (ICALP 2019), and Bandyapadhyay, Fomin & Simonov, (ICALP 2021; JCSS 2024), respectively. Along the way, we obtain faster EPASes for uncapacitated $k$-Median/Means, improving upon the running time of the algorithm by Abbasi et al. (FOCS 2023).
@@ 1 Introduction
Clustering plays an indispensable role in our understanding of very large scale data. In $k$-clustering problems, we are given a set $\mathcal{P}$ of data points in a metric space $\mathcal{M}$, and the goal is to partition $\mathcal{P}$ into $k$ partitions (namely, clusters) such that nearby points should ideally be in the same part, and each cluster is represented by a center. Clustering problems have been studied extensively in theoretical computer science, machine learning, computer vision, among other fields. Among the various notions of clustering, $k$-Median and $k$-Means are arguably the most popular and have been studied algorithmically for decades [KMN+02, HM04a, BCP+24, CEMN22, HM04b, BHI02, KSS10]. Clustering under capacity constraints is a natural and fundamental problem, and hence, quite conceivably, Capacitated $k$-Median and Capacitated $k$-means have garnered lot of attention in recent years; see [BFRS15, CR05, Li17, Li16, DL16, CGTS02, Coh20].
Capacitated $k$-Median and Capacitated $k$-means are inherently harder than the vanilla (Uncapacitated) $k$-Means or $k$-Median, since there is hard constraint on the number of clusters as well as the number of points that can be assigned to each cluster. In fact, till-date the best known approximation for capacitated $k$-median remains $O(\log k)$ (folklore), using the probabilistic embedding of metric spaces into trees (Adamczyk et al. gave an explicit proof in [ABM+19] for completeness). Constant-factor approximation algorithms exist when either the capacities [Li17, Li16], or the number of clusters [DL16, BRU16], are allowed to be exceeded by a $(1+\varepsilon)$ factor, for some constant $\varepsilon>0$. From the lower bounds side, surprisingly, no better lower bound than $(1+\frac{2}{e}-\varepsilon)$ is known, which is inherited from the uncapacited version. Given this evident gap, it is natural to ask whether the approximation landscape can be improved for metric spaces with special properties, possibly at the cost of higher running time. While it remains an open question whether a constant-factor approximation algorithm exists for general metrics that runs in polynomial time, Cohen-Addad and Li [CL19a], in 2019, showed a $(3+\varepsilon)$-approximation for the problem, running in time FPT in $k$ and $\varepsilon$. However, even in the uncapacitated case and with FPT running time, the lower bound remains at $1+\frac{2}{e}$, assuming Gap-ETH. Notably, for uncapacitated case, this lower bound can be matched by an FPT algorithm [CGK+19]. A central open question, thus, is: For which families of metric spaces, is it possible to obtain better upper bounds for capacitated $k$-Median/Means.The only improvement in the affirmative sense is known for Euclidean spaces. For instance, for Euclidean spaces, Cohen-Addad and Li showed that there exists $(1+\varepsilon)$-approximation algorithms for both problems with running time $f(\varepsilon,k)n^{O(1)}$ [CL19a] (see also [BJK18]).  [^1] In this context, a recent work of [HLLW25] on coresets for constrained clustering is also relevant. We discuss the implications of this work and its relation with our work in Section 4.4.
As mentioned before, stronger upper bounds are known for uncapacitated $k$-Median/Means. Indeed, Cohen-Addad et al. [CGK+19] obtained tight $(1+\frac{2}{e}+\varepsilon)$ and $(1+\frac{8}{e}+\varepsilon)$ approximation factors for $k$-Median and $k$-Means in general metric spaces, respectively, with running times FPT in $k$ and $\varepsilon$. However, an important question remains: Can the lower bound barriers for uncapacitated $k$-Median/Means be broken for special metric spaces? To that end, improved FPT-approximation algorithms have been developed for Euclidean spaces; see [BHI02, KSS10]. Bădoiu, Har-Peled and Indyk [BHI02] showed that, for points in $\mathbb{R}^{d}$, a $(1+\varepsilon)$-approximate $k$-Median can be computed in expected $2^{{k/\varepsilon}^{O(1)}}d^{O(1)}n\log^{O(k)}n$ time. Subsequently, Har-Peled and Mazumdar in their major work [HM05], showed the existence of small coresets for the problems of computing $k$-Median and $k$-Means clustering for points in low dimension. Kumar, Sabharwal and Sen in their breakthrough work [KSS10] gave simple randomized algorithms for $k$-Median/Means that obtain $(1+\varepsilon)$-approximation solutions with constant probability, running in time $O(2^{{(k/\varepsilon)}^{O(1)}}dn)$. A fundamental aspect of these works is that they primarily rely on key sampling ideas (or, roughly speaking, coresets) that crucially exploits properties of the Euclidean space. Moreover, all these works primarily focus on the continuous Euclidean cases (where centers can be opened anywhere), except that the result of Kumar et al. [KSS10] also works for discrete $k$-Means. Feldman et al. [FSS20] obtained coresets for $k$-Means, PCA and projective clustering whose sizes are independent of the dimension. Later, Sohler and Woodruff [SW18] generalized these results and obtained coresets for $k$-Median and for subspace approximation of sizes independent of the dimension $d$. Braverman et al. [BJKW21] designed coresets for for $k$-Median in (the shortest-path metric of) an excluded-minor graph, which is also applicable to Euclidean metrics by using terminal embedding result of Narayanan and Nelson [NN19] that extends the Johnson-Lindenstrauss Lemma. After a long sequence of work [CAFS21, HV20, BBH+20, BJKW21, FL11], finally in 2021, Cohen-addad, Saulpic, and Schwiegelshohn [CSS21] gave a unified framwork for a large variety of metrics, ranging from Euclidean space, doubling metric, minor-free metric, as well as capturing the general metric case. However, while these results provide excellent bounds on coreset size, a key concern remains: the computational cost of clustering using them. Hence, a natural question to ask: Can coresets be used to obtain faster algorithms for $k$-Clustering?
@@
Efficient parameterized approximation scheme (EPAS) [^2] A $(1+\varepsilon)$ approximation algorithm that runs in time $f(k,\varepsilon)\operatorname{poly}(n)$ for every $\varepsilon>0$. is perhaps, the best combination of approximation and parameterized algorithms, and has been a majorly popular theme in recent algorithms research (see [ABB+23, BFM24, Mat00, ORSS13, DX20] and the references therein). However, for $k$-Median/Means, unfortunately, it is not possible to obtain EPASes for points in general metric spaces [Das08, ACKS15]. Therefore, quite conceivably, much of the work focused on designing EPASes for points in special metrics, e.g., continuous Euclidean spaces [KSS10, BHI02]. In a recent work [ABB+23], Abbasi et al. presented a unified EPAS which works for various clustering objectives as well as diverse metric spaces. To that end, they introduced the notion of $\varepsilon$-scatter dimension (see Section 2.1 for definition), which includes various commonly studied metric spaces, including, continuous Euclidean space, doubling spaces, etc. Very recently, Bourneuf and Pilipczuk [BP25] showed that metrics induced by graphs from any fixed proper minor-closed graph class also have bounded scatter dimension. Moreover, they introduced metric analogs of well-known graph invariants from the theory of sparsity, and as a consequence of this toolbox showed a coreset for $k$-Center in any proper minor-closed graph class whose size is polynomial in $k$ but the exponent of the polynomial depends on the graph class and $\frac{1}{\varepsilon}$. Abbasi et al. [ABB+23] bypassed coresets to obtain their clustering results. They mentioned that, obtaining clustering algorithms via coresets faces information-theoretic limitation, particularly in the context of $k$-Center, since coresets of desirable sizes do not exist in high-dimensional Euclidean spaces [BJKW19]. However, a seemingly inherent limitation of their approach was that they could only handle clustering objectives, where each point was assigned to the closest chosen center, i.e., Voronoi clustering [^3] A recent work of Gadekar and Inamdar [GI25] also used scatter dimension to design an EPAS for a common generalization of $k$-Center and $k$-Median that is not captured by the general norm framework of [ABB+23]; demonstrating that the notion of (algorithmic) scatter dimension has a wider applicability in clustering. However, the clustering objective considered in [GI25] still can only handle Voronoi assignments.. Indeed, in their work they explicitly asked the following open question:
@@
Is it possible to design EPASes for constrained $(k,z)$-Clustering in metric spaces of bounded (algorithmic) scatter dimension, for assignment constraints such as capacities, fairness, and diversity?
@@
In fact, [ABB+23] asked this question for a more general problem, called constrained $k$-Clustering. However, this problem is so general that it also captures Capacitated $k$-Center, which is known to be W$[1]$-hard to approximate to a factor arbitrarily close to $1$ in metric spaces of bounded highway dimension [FV22]. Furthermore, it has been observed that metric spaces of bounded highway dimension have bounded scatter dimension [^4] See the talk on [ABB+23] at FOCS 2023: https://focs.computer.org/2023/schedule/., implying that the original question of [ABB+23] cannot be answered in the affirmative. Therefore, we reinterpret their question by restricting our focus to constrained $(k,z)$-clustering, where the objective is to minimize the sum of $z$th power of distances. This problem is general enough to capture many well studied clustering problems (e.g., $z=1,2$, correspond to $k$-Median and $k$-Means, respectively), but also manages to bypass the hardness of result of [FV22].

### v3

# Clustering under Constraints: Efficient Parameterized Approximation Schemes

@@ Abstract
Constrained $k$-clustering is a general framework for clustering that captures a wide range of structural and assignment constraints, including capacities, fairness, outliers, and matroid restrictions. Despite its broad applicability, efficient parameterized approximation schemes (EPASes) for these problems were previously known only in continuous Euclidean spaces or for formulations restricted to Voronoi assignments, where each point is assigned to its nearest center. Moreover, it is known that EPASes cannot exist in general metric spaces.
We present a unified framework that yields EPASes for constrained $(k,z)$-clustering in metric spaces of bounded (algorithmic) scatter dimension, a notion introduced by Abbasi et al. (FOCS 2023). They showed that several well known metric families, including continuous Euclidean spaces, bounded doubling spaces, planar metrics, and bounded treewidth metrics, have bounded scatter dimension. Subsequently, Bourneuf and Pilipczuk (SODA 2025) proved that this also holds for metrics induced by graphs from any fixed proper minor closed class. Our result, in particular, addresses a major open question of Abbasi et al., whose approach to $k$-clustering in such metrics was inherently limited to Voronoi-based objectives, where each point is connected only to its nearest chosen center.
As a consequence of our framework, we obtain EPASes for several constrained clustering problems, including capacitated and matroid $(k,z)$-clustering, fault tolerant and fair $(k,z)$-clustering, as well as for metrics of bounded highway dimension. In particular, our results on capacitated and fair $k$-Median and $k$-Means provide the first EPASes for these problems across broad families of structured metrics. Previously, such results were known only in continuous Euclidean spaces, due to the works of Cohen-Addad and Li (ICALP 2019) and Bandyapadhyay, Fomin, and Simonov (ICALP 2021; JCSS 2024), respectively. Along the way, we also obtain faster EPASes for uncapacitated $k$-Median and $k$-Means, improving upon the running time of the algorithm by Abbasi et al. (FOCS 2023).
@@ 1 Introduction
Clustering and facility location problems form a central theme in modern algorithmic theory. Given a set of points with pairwise distances, the goal is to choose a small number of representative centers and assign points to them in a way that minimizes an aggregate measure of distance. This paradigm captures a wide range of applications in data analysis, machine learning, and operations research. Traditionally, clustering is viewed as partitioning the dataset, where each point belongs to a single cluster represented by its nearest center. A broader and more flexible perspective, however, treats these problems as assignment optimization: each point can distribute its weight among multiple facilities under certain structural or resource constraints (see, e.g.,  [STA97, CG05, KT02, BBG09]).
Among the best known formulations of clustering are the classical $k$-Median and $k$-Means problems, which have been studied intensively across theoretical computer science, machine learning, and optimization for several decades (see, e.g.,  [KMN+02, HM04, BHI02, CEM+22, KSS10]). These models seek $k$ centers that minimize the total (or squared) distance from each point to its assigned center. Over time, several more realistic and constrained variants have been explored. Among these, the Capacitated $k$-Median and $k$-Means problems stand out, where each center can serve only a limited number of points or must satisfy additional combinatorial restrictions on the set of open centers. Such constrained variants naturally arise in network design, resource allocation, and facility planning, and are substantially harder from an algorithmic standpoint. Despite decades of work, Capacitated $k$-Median remains one of the central open problems in approximation algorithms: the best known approximation ratio in general metrics is $O(\log k)$, obtained via probabilistic embeddings of metrics into trees [ABM+19]. Constant factor approximations are known only when either the capacities [LI17, LI16] or the number of clusters [DL16, BRU16] are allowed to be violated by a $(1+\varepsilon)$ factor. On the hardness side, the strongest lower bound known, $(1+\frac{2}{e}-\varepsilon)$ under Gap-ETH, is inherited directly from the uncapacitated version. These gaps between upper and lower bounds have motivated the search for improved approximations in structured metrics, often by allowing fixed parameter or quasi polynomial running time [CL19, BJK18].
In recent years, the connection between the geometric structure of metrics and the design of parameterized approximation algorithms has become increasingly prominent. Cohen-Addad and Li [CL19] obtained a $(3+\varepsilon)$ approximation for Capacitated $k$-Median in general metrics with running time fixed parameter tractable (FPT) in $k$ and $\varepsilon$. Notably, for the uncapacitated case, the lower bound of $(1+\frac{2}{e})$ can be matched by an FPT algorithm [CGK+19b]. A central question thus arises: for which families of metric spaces can one obtain better upper bounds for Capacitated $k$-Median or $k$ Means? The only affirmative progress is known for Euclidean metrics, where Cohen-Addad and Li showed that $(1+\varepsilon)$ approximation algorithms exist for both problems with running time $f(\varepsilon,k)\,n^{O(1)}$ [CL19] (see also [BJK18]). [^1] In this context, a recent work of Huang et al. [HLL+25] on coresets for constrained clustering is also relevant; we discuss its implications in Section 4.3.
As mentioned before, stronger upper bounds are known for the uncapacitated versions of the problem. Indeed, Cohen-Addad et al. [CGK+19b] obtained tight $(1+\frac{2}{e}+\varepsilon)$ and $(1+\frac{8}{e}+\varepsilon)$ approximation factors for $k$-Median and $k$-Means, respectively, in general metric spaces, with running times fixed parameter tractable (FPT) in $k$ and $\varepsilon$. However, an important question remains: can these lower bound barriers for uncapacitated $k$-Median and $k$-Means be broken for structured metric spaces? This question has driven a series of advances in Euclidean settings, where geometric structure enables substantially stronger approximation guarantees. Bădoiu, Har-Peled, and Indyk [BHI02] showed that, for points in $\mathbb{R}^{d}$, a $(1+\varepsilon)$-approximate $k$-Median can be computed in expected time $2^{(k/\varepsilon)^{O(1)}}d^{O(1)}n\log^{O(k)}n$. Har-Peled and Mazumdar [HM05] subsequently proved the existence of small coresets for both $k$-Median and $k$-Means in low dimensions, while Kumar, Sabharwal, and Sen [KSS10] gave simple randomized algorithms achieving $(1+\varepsilon)$ approximations with constant probability in $O(2^{(k/\varepsilon)^{O(1)}}dn)$ time. A common theme among these works is the use of geometric sampling and coreset constructions that exploit Euclidean structure. Most of these algorithms address the continuous setting, where centers may be placed anywhere in $\mathbb{R}^{d}$, although the result of [KSS10] also applies to the discrete $k$-Means problem.
Over the past decade, coreset techniques have evolved into a general algorithmic framework for clustering across diverse metric spaces. Feldman et al. [FSS20] constructed coresets for $k$-Means, PCA, and projective clustering whose sizes are independent of the dimension, and Sohler and Woodruff [SW18] extended these ideas to $k$-Median and subspace approximation. Braverman et al. [BJK+21] developed coresets for $k$-Median in excluded-minor metrics. After a long sequence of advances [FL11, BBH+20, HV20, BJK+21], Cohen-Addad et al. [CSS21] presented a unified coreset framework encompassing Euclidean, doubling, and minor-free metrics, as well as general metric spaces. While these results provide remarkably small coresets and broad applicability, an important algorithmic challenge remains: Can coreset-based methods also yield unified framework for designing faster (EPASes) approximation schemes for (constrained) Clustering?
@@
An efficient parameterized approximation scheme (EPAS) [^2] A $(1+\varepsilon)$-approximation algorithm that runs in time $f(k,\varepsilon)\operatorname{poly}(n)$ for every $\varepsilon>0$. is perhaps the most powerful synthesis of approximation and parameterized algorithms, and has become a central theme in recent algorithmic research (see [ABB+23, BFM24b, MAT00, ORS+13, DX20] and the references therein). However, for $k$-Median and $k$-Means, it is impossible to obtain EPASes for arbitrary metric spaces [CGK+19a] under standard complexity-theoretic conjectures. Consequently, much of the literature has focused on designing EPASes for special metrics, such as continuous Euclidean spaces [KSS10, BHI02].
In a recent work, Abbasi et al. [ABB+23] presented a unified EPAS framework that applies to a variety of clustering objectives as well as diverse metric spaces. To achieve this, they introduced the notion of $\varepsilon$-scatter dimension (see Section 2.1 for the formal definition), which encompasses many commonly studied metric families, including continuous Euclidean and doubling spaces. Subsequently, Bourneuf and Pilipczuk [BP25] showed that metrics induced by graphs from any fixed proper minor-closed graph class also have bounded scatter dimension. Moreover, they developed metric analogs of well-known graph invariants from the theory of sparsity, and as a consequence constructed a coreset for $k$-Center in any proper minor-closed graph class whose size is polynomial in $k$, with the exponent depending on the graph class and $\frac{1}{\varepsilon}$.
Abbasi et al. [ABB+23] bypassed the use of coresets to obtain their clustering results. They argued that deriving clustering algorithms via coresets faces information-theoretic limitations, particularly for $k$-Center clustering, since small coresets do not exist in high-dimensional Euclidean spaces [BJK+19]. However, an inherent limitation of their approach is that it only applies to clustering objectives in which each point is assigned to its nearest chosen center, i.e., Voronoi clustering. [^3] A recent work of Gadekar and Inamdar [GI25] also used scatter dimension to design an EPAS for a common generalization of $k$-Center and $k$-Median that is not captured by the general norm framework of [ABB+23], further demonstrating the broader applicability of the scatter-dimension concept. However, the clustering objective considered in [GI25] still only supports Voronoi assignments. Indeed, Abbasi et al. explicitly posed the following broad open question:
@@
Is it possible to design EPASes for constrained $(k,z)$-Clustering [^4] Informally, the objective of the $(k,z)$-clustering includes the sum of $z$-th power of distances, where $1\leq z<\infty$ is fixed. For example, $z=1,2$ corresponds to $k$-Median and $k$-Means, respectively. A formal definition follows. in metric spaces of bounded (algorithmic) scatter dimension, for assignment constraints such as capacities, fairness, and diversity?
@@
In fact, [ABB+23] asked this question for an even more general constrained clustering problem; however that problem is so general that it also captures the Capacitated $k$-Center problem, which is known to be W[1]-hard to approximate within any factor arbitrarily close to $1$ in metric spaces of bounded highway dimension [FV22]. Furthermore, it has been observed that metric spaces of bounded highway dimension also have bounded scatter dimension;  [^5] See the talk on [ABB+23] at FOCS 2023: https://focs.computer.org/2023/schedule/. implying that the original question of [ABB+23] cannot be answered in the affirmative. Therefore, we reinterpret their question by restricting our focus to constrained $(k,z)$-Clustering, while still avoiding the hardness barrier of [FV22].
The above discussion motivates a more general view of clustering under constraints. Rather than focusing on any specific objective or restriction, we consider the broad setting of constrained $(k,z)$-clustering, which captures a wide spectrum of problems including capacitated, fault-tolerant, fair, outlier, and matroid-constrained variants. Our goal is to develop a unified algorithmic framework that yields efficient parameterized approximation schemes for this family of problems in metric spaces of bounded (algorithmic) scatter dimension, thereby also addressing the underlying challenge posed in Question 1. Our formulation is deliberately generic: it views clustering as an assignment optimization problem, where each data point may fractionally distribute its weight among multiple centers under feasibility constraints. This unified perspective underlies our results and enables reasoning about diverse constrained objectives within a single framework.
We now formally define the problem.
1.1 Problem Formulation
We consider a generalized notion of $(k,z)$-Clustering, which we call constrained $(k,z)$-Clustering. The input consists of a set of points $P$ (also called clients) with $|P|=n$, a set of facilities $F$, and a metric space $M=(P\cup F,d)$. Each point $p\in P$ has a nonnegative weight $w(p)$, and we are given a positive integer $k$.
The goal is to choose a center set $X\subseteq F$ with $|X|=k$, together with an assignment function $f:P\times X\rightarrow\mathbb{R}_{\geq 0}$ satisfying
$\sum_{x\in X}f(p,x)=w(p)\qquad\text{for all }p\in P,$
so as to minimize the clustering cost
${\sf cost}(X)=\sum_{p\in P}\sum_{x\in X}f(p,x)\cdot d(p,x)^{z},$
for a fixed constant $z\geq 1$. In addition, we impose two types of constraints:
(i) center constraints
that restrict the allowable center sets $X$, and
(ii) assignment constraints
that restrict feasible assignments $f$ for a given feasible center set.
This general framework captures a wide range of well-studied constrained clustering problems. Below, we describe several important special cases.
  • Capacitated Clustering [LI17, DL16, KS00, CHK12, CL19] In this setting, each facility $c\in F$ has a capacity $\eta_{c}$ such that the total points assigned to $c$ are at most $\eta_{c}$. This is captured by the constraint for all $c\in C$ that $\sum_{p\in P}f(p,x)\leq\eta_{c}$, for all $p\in P$. [^6] We consider a general setting of capacitated clustering, similar to [CL19], where points can be assigned fractionally to multiple facilities.
  • Fault Tolerant Clustering [HHL+16, KPS00, BSS10]: In this problem, each point $p\in P$ is assigned to $\ell$ closest centers in $X$ to $p$. This can be obtained by the constraint $f(p,x)\leq\tfrac{w(p)}{\ell}$ for all $p\in P$ and $x\in X$, and considering the normalized cost $\sum_{p\in P}\sum_{x\in X}f(p,x)\cdot(d(p,x))^{z}$, which is $(\tfrac{1}{\ell})^{th}$ factor of the original cost of $X$.
  • $(\alpha,\beta)$-Fair Clustering [CKL+17, BFS24a]: This problem partitions $P$ into $\ell$ colors $P_{1},\dots,P_{\ell}$, and we are given a range $\alpha_{i},\beta_{i}]$ for each color $i\in[\ell]$. The goal is to find a set $X$ of $k$ centers and a clustering of $P$ such that, in each cluster, the proportion of color-$i$ points to the cluster points is in the range $\alpha_{i},\beta_{i}]$. This problem is captured by adding the following constraint on $f$: for each center $x\in X$ and each color $i\in[\ell]$, we have $\frac{\sum_{p\in P_{i}}f(p,x)}{\sum_{p\in P}f(p,x)}\in[\alpha_{i},\beta_{i}]$.
  • Outlier Constraint [CKM+01, CHE08]: Here, we are given additional parameter $M\leq W$, where $W=\sum_{p}w(p)$, and the constraint is to assign all the weight of points except $M$ to $X$. This can be achieved by relaxing the constraint $\sum_{x\in X}f(p,x)=w(p)$ to $\sum_{x\in X}f(p,x)\leq w(p)$, and adding the constraint $\sum_{p\in P}\sum_{x\in X}f(p,x)\geq W-M$.
  • Matroid constrained Clustering [KLS18, KKN+11, SWA16]: In this problem, we are additionally given a matroid $\mathbb{M}$ over $F$, and the constraint is to have a center set $X$ that is an independence set in $\mathbb{M}$. A particular example of this problem is a fair variant of $k$-median [TGO+22], where each facility has one of the $k$ colors, and there is a bound $\alpha_{i}$ for each color. The constraint is to choose $X$ such that it has exactly $\alpha_{i}$ centers of color $i$.
1.2 Scatter Dimension vs. Algorithmic Scatter Dimension
Our algorithms work for metric spaces with bounded scatter dimension. At a high level, the scatter dimension [^7] For simplicity, we focus on a particular metric space. These definitions can analogously be extended to a class of metric spaces; see Section 2. of a metric space $M=(P\cup F,d)$ is the length of the longest sequence of center–point pairs $(x_{1},p_{1}),(x_{2},p_{2}),\dots$, where $(x_{i},p_{i})\in F\times P$ such that (i) $d(x_{i},p_{i})>1+\varepsilon$ for all pairs in the sequence, and (ii) $d(x_{j},p_{i})\leq 1$ for all $i<j$. Thus, this is a combinatorial property of a metric space that bounds the length of such sequences, which are called $\varepsilon$-scattering sequences. Note that this notion neither relies on nor requires the existence of an algorithm to compute such a sequence.
On the other hand, the algorithmic scatter dimension of $M$ is the length of the longest $\varepsilon$-scattering sequence computed by the best ball-intersection algorithm, where “best” refers to an optimized procedure that produces the smallest possible length of such sequence for arbitrary choices of points. However, to design approximation algorithms under bounded scatter dimension, one still requires the existence of a ball-intersection algorithm for the metric space. In fact, for our framework, we require a stronger ball-intersection algorithm, namely $\mathcal{A}_{B}$ mentioned above, which not only satisfies the distance constraints but also returns a feasible center.